5 consecutive even integars whose sum is -180.
So far I have:
-18*10
-10*18
Are these right?

Answer by edjones(8007)

Let 5 consecutive even integers be x, x+2, x+4, x+6, x+8
The formula for 5 consecutive odd integers would be the same.
x+x+2+x+4+x+6+x+8=180
5x+20=180
5x=160
x=32
Check:
32+34+36+38+40=180 true
Ed